Mother spent 1/2 of her money in the first store. She spent 2/3 of her money in the second store. She has $25 left. How much money does she have first?
Answer by checkley77(12569)

x=25+2/3*1/2x+1/2x
x=25+2x/6+x/2
x=25+(2x+3x)/6
x=25+5x/6
x-5x/6=25
x/6=25
x=25*6
x=$150 initial amount Mother had.
Proof:
150=25+2/3*150/2+150/2
150=25+300/6+150/2
150=25+50+75
150=150